. A binary operation is defined as p*q = q-3p, evaluate:
a. 2*1
b. 1*2
c. Is the operation * commutative?
d. (1*2)*3
Show solution

Answer by ikleyn(52781)

.

I will give you an idea "how to start"

(a)  2*1.

     Substitute p= 2, q= 1 into the formula


        2*1 = 1 - 3*2 = 1 - 6 = -5.



(b)  Do (b) exactly in the same way.



(c)  Now, when you just know how to handle two-term aggregate,

     move forward and calculate (1*2) first.


     After that, make the last step.

Having my instructions, you just know what to do and how move forward.

So, I hope you can complete the rest on your own.


If you have questions, let me know.